# Process-Supervised LLM Recommenders via Flow-guided Tuning

### v1

@@ Abstract.
While large language models (LLMs) are increasingly adapted for recommendation systems via supervised fine-tuning (SFT), this approach amplifies popularity bias due to its likelihood maximization objective, compromising recommendation diversity and fairness. To address this, we present Flow-guided fine-tuning recommender (Flower), which replaces SFT with a Generative Flow Network (GFlowNet) (Bengio et al., 2024b) framework that enacts process supervision through token-level reward propagation. Flower’s key innovation lies in decomposing item-level rewards into constituent token rewards, enabling direct alignment between token generation probabilities and their reward signals. This mechanism achieves three critical advancements: (1) popularity bias mitigation and fairness enhancement through empirical distribution matching, (2) preservation of diversity through GFlowNet’s proportional sampling, and (3) flexible integration of personalized preferences via adaptable token rewards. Experiments demonstrate Flower’s superior distribution-fitting capability and its significant advantages over traditional SFT in terms of fairness, diversity, and accuracy, highlighting its potential to improve LLM-based recommendation systems. The implementation is available via https://github.com/Mr-Peach0301/Flower.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
In recent years, recommendation systems powered by Large Language Models (LLMs) have emerged as a rapidly advancing field of research (Wu et al., 2024b; Lin et al., 2024; Harte et al., 2023a). Leveraging LLMs’ vast repository of world knowledge and advanced learning capabilities, LLMs have demonstrated remarkable potential to enhance recommendation accuracy (Bao et al., 2024a; Zhang et al., 2023a), improve fairness (Dai et al., 2024; Jiang et al., 2024; Deldjoo and Di Noia, 2024), and deliver more explainable (Yang et al., 2024b; Ma et al., 2024) and controllable (Chen et al., 2025; Lu et al., 2024; Shen et al., 2024) recommendation results. These advancements contribute significantly to improving the overall user experience in recommendation systems.
@@
While effective in general-purpose tasks, pretrained LLMs require supervised fine-tuning (SFT) to acquire knowledge specific to downstream tasks. In recommendation tasks, preference modeling is a key objective. A commonly used SFT paradigm involves constructing instruction data from historical user behavior sequences, enabling the LLM to predict the next item based on prior interactions (Bao et al., 2023a, 2024b; Zhang et al., 2024c; Gao et al., 2024). Items are typically represented by their titles, which consist of multiple tokens generated sequentially by the LLM. SFT leverages the cross-entropy (CE) loss to maximize the likelihood of the target labels. For example, in a next-movie recommendation task (Fig. 1(a)), the movie title “Back to the Future” is generated token by token. Given the prompt and the first two tokens, “Back” and “to”, SFT optimizes the prediction of the next ground-truth token, “the”. While effective, SFT introduces two critical challenges in recommendation systems:
  • Limited diversity in recommendations. SFT often encourages models to generate more “conventional” outputs, prioritizing high-probability tokens and sacrificing diversity and creativity (Li et al., 2024). This tendency can lead to overfitting dominant patterns in the fine-tuning dataset, resulting in homogeneous recommendations that lack personalization, which undermines the core objectives of recommendation systems.
@@
  • Amplification of popularity bias. SFT tends to reinforce pre-existing biases in both the fine-tuning dataset and the pre-training corpora (Dai et al., 2024). This can cause models to over-recommend popular items, reducing exposure to less popular or niche content, and ultimately degrading user experience and the fairness of recommendations.
@@
To address these limitations, existing solutions in LLM-based recommendation systems (LRSs) can be broadly categorized into two approaches. The first approach involves modifying the SFT learning process. For example, assigning different weights to samples to adjust the learning loss (Jiang et al., 2024) can mitigate category-specific popularity bias but fails to enhance the diversity of LLM-generated outputs. Another strategy is multi-stage SFT, where expert priors are integrated into successive stages to promote diversity (Chen et al., 2025). However, this method relies heavily on manually designed, complex multi-stage workflows, which are prone to error propagation and may exacerbate bias. The second approach emphasizes post-SFT policy optimization to better align LLMs with human preferences, aiming to alleviate popularity bias and enhance diversity. For example, reinforcement learning with human feedback (RLHF) leverages reward signals to guide the recommender toward diverse outputs (Lu et al., 2024). However, RLHF’s reward maximization objective often leads to a collapse toward high-reward outcomes, which can further reduce diversity (Kirk et al., 2024; Yu et al., 2024). More recently, Direct Preference Optimization (DPO) (Rafailov et al., 2024) has been applied to LRSs, using contrastive positive and negative samples to reduce popularity bias (Liao et al., 2024a). Despite its promise, DPO tends to induce distributional collapse by driving the probabilities of negative samples to zero (Azar et al., 2024), ultimately limiting recommendation diversity (Gao et al., 2024).
@@
By leveraging GFlowNet’s mechanism, we conceptualize the prefix tree of all feasible items as an irreversible flow network, where “flow” refers to the unnormalized probabilities moving from the root node to the leaf nodes. The reward assigned to each item corresponds to the flow at the leaf nodes. This framework enables us to compute the flow values at each branching point of the network, which serve as token-level rewards for the next-token prediction task. With these token-level rewards, Flower supervises the generation probabilities of each token during fine-tuning. As shown in Fig. 1(b), unlike SFT, which focuses on optimizing the prediction of a single ground truth token at a time, the objective of Flower is to align the policy (i.e., the predicted probability distribution over the token) with the token-level reward distribution across all feasible tokens. Beyond fairness and diversity considerations, we further incorporate personalized user preferences into the derived token-level rewards using an auxiliary model, thereby enhancing both the accuracy and personalization of the recommendations.
Flower is a fine-tuning paradigm designed to enhance performance over conventional SFT. After tuning with Flower, additional alignment methods such as RLHF and DPO can still be applied. A comparison of different methods is provided in Table 1. The main contributions of this paper are as follows:
@@ 2. Preliminary
In this section, we present SFT for the LLM-based next-item recommendation task, along with an introduction to the GFlowNets framework and its training objectives.
@@ 2.1. SFT for Next-item Recommendation
Next-item prediction is a core task in recommender systems. This task requires the model to predict the next item a user is likely to engage with, given a sequence of historically interacted items (i.e., behavior sequence) as input. Leveraging the powerful generative capabilities of LLMs, this task can be formulated as directly generating the next recommended item. Following the BIGRec model (Bao et al., 2023a), we organize the training process in the form of instruction tuning. The task can be formally defined as follows:
Given a prompt $\mathbf{x}$, which typically includes a description of the task and a sequence of historically interacted items for a user, the policy $\pi_{\theta}$ is trained to generate the next item $\mathbf{y}$. Here, the next item $\mathbf{y}$ is represented by its title, consisting of multiple tokens: $\mathbf{y}=y_{1:T}=[y_{1},y_{2},\cdots,y_{T}]$.
During SFT, The cross-entropy loss used in teacher forcing of sequence-to-sequence models can be expressed in detail as follows:
(1) $\mathcal{L}_{\text{SFT}}=-\frac{1}{T}\sum_{t=1}^{T}\log\pi_{\theta}(y_{t}\mid\mathbf{x},y_{1:t-1}).$
where $y_{t}$ is the correct token, $\pi_{\theta}(y_{t}\mid y_{1:t-1},\mathbf{x})$ is the probability predicted by the model for token $y_{t}$, conditioned on the input sequence $\mathbf{x}$ and the previously generated tokens $y_{1:t-1}$.
After fine-tuning, the LLM is expected to generate items as recommendations during the inference stage. However, LLMs may occasionally generate nonexistent items. To address this, Bao et al. (2023a) propose a matching mechanism to map generated items to real ones, while Bao et al. (2024b) introduce constrained decoding, limiting token sampling to ensure valid item generation. We adopt the latter approach to ensure recommendation validity.
@@ 2.2. Problems in SFT-based Recommendations
SFT introduces bias and fairness issues in LRSs (Dai et al., 2024; Jiang et al., 2024; Gao et al., 2024). To illustrate this, we utilize the CDs and Video datasets from the Amazon Review datasets, dividing all items into five groups based on their frequency of occurrence in the historical sequences of the training data. After fine-tuning the LLM with BIGRec [^1] Detailed experimental settings are provided in Section 4, we analyze the distribution of the top-5 recommended items across these five groups. Furthermore, to investigate the influence of the temperature parameter during the inference stage, we conduct experiments with temperatures set to 1.0, 1.2, 1.5, and 2.0.
@@ 2.3. Generative Flow Networks (GFlowNets)
Generative Flow Networks (GFlowNets) (Bengio et al., 2024a) are a family of probabilistic generative models designed to learn stochastic policies for sequential decision-making. Inspired by physical flow systems, GFlowNets introduce a flow-based framework, where the term “flow” refers to the underlying unnormalized probabilities that can be learned and utilized to model diverse outcomes.
At their core, GFlowNets operate on a directed acyclic graph (DAG), where each path represents a sequence of actions leading from a root state to a terminal state. The probability of a sequence is defined by the flow passing through its path. This framework enables the allocation of probability mass proportional to the reward of an outcome, making GFlowNets particularly effective for generating diverse and high-reward solutions.
Formally, let $G=(\mathcal{S},\mathcal{A})$ be a DAG. The vertices $s\in\mathcal{S}$ are referred to as states, and the directed edges $(s_{1}\to s_{2})\in\mathcal{A}$ are called actions. If $(s_{1}\to s_{2})$ is an edge, $s_{2}$ is a child of $s_{1}$, and $s_{1}$ is a parent of $s_{2}$. The graph includes a unique initial state $s_{0}\in\mathcal{S}$ with no parents, while states with no children are termed terminal and collectively denoted as $\mathcal{Y}$. A trajectory or action sequence $\tau\in\mathcal{T}$ is defined as a sequence of states $\tau=(s_{m}\to s_{m+1}\to\dots\to s_{n})$, where each $(s_{i}\to s_{i+1})$ is an action.
A (forward) policy is a collection of distributions $P_{F}(-|s)$ over the children of each nonterminal state $s\in\mathcal{S}$. This forward policy induces a distribution over trajectories $\mathcal{T}$ as follows:
(2) $P_{F}(\tau=(s_{0}\to\dots\to s_{n}))=\prod_{i=0}^{n-1}P_{F}(s_{i+1}|s_{i}).$
Note that any distribution over trajectories that arises from a policy satisfies a Markov property: the marginal choice of action $(s_{i}\to s_{i+1})$ out of a state $s_{i}$ is independent of how $s_{i}$ was reached.
@@
2.4. Training Objectives for GFlowNets
The sum over all $\tau\in\mathcal{T}$ in Eq. (3) is intractable to compute directly, necessitating the use of auxiliary variables to facilitate learning the policy. These auxiliary variables represent different formulations of flow, such as the edge flow $F(s\to t)$ used in Flow Matching (Bengio et al., 2024a), state flow $F(s)$ used in Detailed Balance (DB) (Bengio et al., 2024b; Malkin et al., 2024), and initial state flow $Z$ used in Trajectory Balance (TB) (Malkin et al., 2024). In this work, we focus on the Subtrajectory Balance (SubTB) objective (Madan et al., 2023), which generalizes the DB and TB conditions as special cases.
Subtrajectory Balance. Let $F(s)$ be a state flow function. For any trajectory segment $\tau_{m:n}=(s_{m}\to\cdots\to s_{n})$, the following constraint must hold:
@@ 2.4. Training Objectives for GFlowNets
where $F(y)=R(y)$ if $y$ is terminal, and $F(s_{0})=Z$ if $s_{0}$ is the initial state. Here, $P_{F}(s_{i+1}|s_{i};\theta)$ is the forward policy, parameterized by $\theta$, representing the probability of taking action $(s_{i}\to s_{i+1})$ conditioned on the state $s_{i}$. Similarly, $P_{B}(s_{i}|s_{i+1};\theta)$ is the backward policy, modeling the reverse transitions.
It is worth noting that the DB constraint (Bengio et al., 2024b) is a special case of Eq. (4) when the trajectory segment consists of a single action, i.e., $\tau=(s_{m},s_{m+1})$. Likewise, the TB constraint (Malkin et al., 2024) corresponds to the case where $\tau$ spans the entire trajectory, i.e., $\tau=(s_{0},s_{1},\cdots,y)$.
@@ 3.1. Problem Formulation
Let $\mathcal{I}$ denote the set of all valid items in the data set. Each item $i\in\mathcal{I}$ is represented by its title, a token sequence $\mathbf{y}^{(i)}=y_{1:T}^{(i)}=[y_{1}^{(i)},y_{2}^{(i)},\cdots,y_{T}^{(i)}]$, where $y_{t}^{(i)}$ is a token in the token space of the LLM. The titles of all items collectively form a prefix tree, where items share common prefixes. For example, the movie titles “Back to the Future II” and “Back to Life” share the prefix “Back to” but diverge at subsequent tokens into two branches: “the Future II” and “Life”. Note that in this example, words are used as tokens for illustration purposes. In practice, an item is automatically tokenized by an open-source LLM, such as LLaMA or Qwen.
The prefix tree of all items forms a special DAG, where each node represents an exact token, and each directed edge corresponds to a possible next token. The root node represents the empty sequence, while paths from the root to terminal nodes represent complete item titles. The state and action are defined as follows:
  • The state $s_{t}$ for a node $y_{t}$ is defined as a partial token sequence from the root node to this node, i.e., a prefix of one or more item titles: $s_{t}=y_{\leq t}=[y_{0},\cdots,y_{t}]$.
  • An action $a_{t}$ is defined as the transition from the current state $s_{t}$ to a child state $s_{t+1}$ by appending a valid next token $y_{t+1}$ to the current sequence.
An example of this tree structure for all 7 movies in Fig. 3(a) is illustrated in Fig. 3(b). At the node “to”, the state is “Back to”, and three possible actions include transitioning to one of the following states: “Back to School”, “Back to Life”, or “Back to the”. The flow through the prefix tree aligns with the GFlowNets framework, where the probability of generating a specific item is distributed along its corresponding path. For instance, the probability of generating the item “Back to Life” is computed as the product of probabilities along the path “Back” $\to$ “to” $\to$ “Life”.
Remark: This formulation enables the model to seamlessly handle sequential generation tasks by interpreting the generation process as a traversal through the prefix tree. The tree structure enforces that only valid tokens are sampled at each step, thereby preventing the generation of invalid items. Additionally, this approach ensures alignment between the model’s outputs and the dataset, leveraging the GFlowNets framework to guarantee diversity and proportionality in the generated recommendations.

### v2

@@ Abstract.
While large language models (LLMs) are increasingly adapted for recommendation systems via supervised fine-tuning (SFT), this approach amplifies popularity bias due to its likelihood maximization objective, compromising recommendation diversity and fairness. To address this, we present Flow-guided fine-tuning recommender (Flower), which replaces SFT with a Generative Flow Network (GFlowNet) (Bengio et al., 2024b) framework that enacts process supervision through token-level reward propagation. Flower’s key innovation lies in decomposing item-level rewards into constituent token rewards, enabling direct alignment between token generation probabilities and their reward signals. This mechanism achieves three critical advancements: (1) popularity bias mitigation and fairness enhancement through empirical distribution matching, (2) preservation of diversity through GFlowNet’s proportional sampling, and (3) flexible integration of personalized preferences via adaptable token rewards. Experiments demonstrate Flower’s superior distribution-fitting capability and its significant advantages over traditional SFT in terms of accuracy, fairness, and diversity, highlighting its potential to improve LLM-based recommendation systems. The implementation is available via https://github.com/Mr-Peach0301/Flower.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729981isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systems
@@ 1. Introduction
In recent years, recommendation systems powered by Large Language Models (LLMs) have emerged as a rapidly advancing field of research (Wu et al., 2024; Lin et al., 2024a). Leveraging LLMs’ vast repository of world knowledge and advanced learning capabilities, LLMs have demonstrated remarkable potential to enhance recommendation accuracy (Bao et al., 2024; Lin et al., 2024b), improve fairness (Dai et al., 2024; Jiang et al., 2024), and deliver more explainable (Yang et al., 2024; Ma et al., 2024) and controllable (Chen et al., 2025; Lu et al., 2024; Shen et al., 2024) recommendation results. These advancements contribute significantly to improving the overall user experience in recommendation systems.
@@
While effective in general-purpose tasks, pretrained LLMs require supervised fine-tuning (SFT) to acquire the knowledge specific to downstream tasks. In recommendation tasks, preference modeling is a key objective. A commonly used SFT paradigm involves constructing instruction data from historical user behavior sequences, enabling the LLM to predict the next item based on prior interactions (Bao et al., 2025, 2024; Gao et al., 2025). Items are typically represented by their titles, which consist of multiple tokens generated sequentially by the LLM. SFT leverages the cross-entropy (CE) loss to maximize the likelihood of the target labels. For example, in a next-movie recommendation task (Fig. 1(a)), the movie title “Back to the Future” is generated token by token. Given the prompt and the first two tokens, “Back” and “to”, SFT optimizes the prediction of the next ground-truth token, “the”. While effective, SFT introduces two critical challenges for recommendation systems:
  • Limited diversity. SFT often drives models to produce high-probability tokens, leading to overfitting on dominant patterns in the training data (Li et al., 2024). As a result, the model generates homogeneous, less personalized recommendations–undermining the core goal of recommender systems.
@@
  • Popularity bias amplification. SFT reinforces biases present in both fine-tuning data and pretraining corpora (Dai et al., 2024), causing models to over-recommend popular items while underexposing niche content, which harms fairness and user experience.
@@
To address these limitations, existing solutions in LLM-based recommendation systems (LRSs) can be broadly categorized into two approaches. The first approach involves modifying the SFT learning process. For example, assigning different weights to samples to adjust the learning loss (Jiang et al., 2024) can mitigate category-specific popularity bias but fails to enhance the diversity of LLM-generated outputs. Another strategy is multi-stage SFT, where expert priors are integrated into successive stages to promote diversity (Chen et al., 2025). However, this method relies heavily on manually designed, multi-stage workflows, which are prone to error propagation and may exacerbate bias. The second approach emphasizes post-SFT policy optimization to better align LLMs with human preferences, aiming to alleviate popularity bias and enhance diversity. For example, reinforcement learning with human feedback (RLHF) leverages reward signals to guide the recommender toward diverse outputs (Lu et al., 2024). However, RLHF’s reward maximization objective often leads to a collapse toward high-reward outcomes, which can further reduce diversity (Kirk et al., 2024; Yu et al., 2024). More recently, Direct Preference Optimization (DPO) (Rafailov et al., 2023) has been applied to LRSs, using contrastive positive and negative samples to reduce popularity bias (Gao et al., 2025; Liao et al., 2024). Despite its promise, DPO tends to induce distributional collapse by driving the probabilities of negative samples to zero (Azar et al., 2024), ultimately limiting recommendation diversity (Gao et al., 2025).
@@
By leveraging GFlowNet’s mechanism, we conceptualize the prefix tree of all feasible items as an irreversible flow network, where “flow” refers to the unnormalized probabilities moving from the root node to the leaf nodes. The reward assigned to each item corresponds to the flow at the leaf nodes. This framework enables us to compute the flow values at each branching point of the network, which serve as token-level rewards for next-token prediction. With these token-level rewards, Flower supervises the generation probabilities of each token during fine-tuning. As shown in Fig. 1(b), unlike SFT, which focuses on optimizing the prediction of a single ground truth token at a time, the objective of Flower is to align the policy (i.e., the predicted probability distribution over tokens) with the token-level reward distribution across all feasible tokens. Beyond fairness and diversity considerations, we further incorporate personalized user preferences into the derived token-level rewards using an auxiliary model, thereby enhancing both the accuracy and personalization of the recommendations.
Flower is a new fine-tuning paradigm superior to conventional SFT in enhancing diversity while maintaining accuracy. After tuning with Flower, additional alignment methods such as RLHF and DPO can still be applied. A comparison of different methods is provided in Table 1. The main contributions are as follows:
@@ 2. Preliminary
This section introduces SFT for next-item recommendation with LLMs, followed by an overview of the basics of GFlowNets.
@@ 2.1. SFT for Next-item Recommendation
Next-item prediction is a fundamental task in recommender systems, where the goal is to predict the next item a user is likely to interact with, based on their historical behavior sequence. Leveraging the generative capabilities of LLMs, this task can be cast as generating the next recommended item directly.
Following the instruction-tuning paradigm of BIGRec (Bao et al., 2025), we define the task as: given a prompt $\mathbf{x}$—typically describing the task and listing the user’s previously interacted items—the model policy $\pi_{\theta}$ is trained to generate the next item $\mathbf{y}$, represented as a token sequence $\mathbf{y}=[y_{1},y_{2},\cdots,y_{T}]$.
The model is fine-tuned using standard cross-entropy loss with teacher forcing:
(1) $\mathcal{L}_{\text{SFT}}=-\frac{1}{T}\sum_{t=1}^{T}\log\pi_{\theta}(y_{t}\mid\mathbf{x},y_{1:t-1}),$
where $\pi_{\theta}$ denotes the model’s predicted probability for token $y_{t}$ given prior tokens and the input prompt.
After fine-tuning, the LLM is used to generate recommendations during inference. However, it may produce invalid or nonexistent items. To mitigate this, Bao et al. (2025) propose a matching mechanism to align outputs with real items, while Bao et al. (2024) introduce constrained decoding to restrict token sampling to valid continuations. We adopt the latter to ensure recommendation validity.
@@ 2.2. Problems in SFT-based Recommendations
SFT introduces bias and fairness issues in LRSs (Dai et al., 2024; Jiang et al., 2024; Gao et al., 2025). To illustrate this, we utilize the CDs and Video datasets from the Amazon Review datasets, dividing all items into five groups based on their frequency of occurrence in the historical sequences of the training data. After fine-tuning the LLM with BIGRec [^1] Detailed experimental settings are provided in Section 4, we analyze the distribution of the top-5 recommended items across these five groups. Furthermore, to investigate the influence of the temperature parameter during the inference stage, we conduct experiments with temperatures set to 1.0, 1.2, 1.5, and 2.0.
@@ 2.3. Generative Flow Networks (GFlowNets)
Generative Flow Networks (GFlowNets) (Bengio et al., 2024a) are a class of generative models that learn stochastic policies for sequential decision-making. Inspired by physical flow systems, GFlowNets define unnormalized probabilistic flows to model diverse outcomes, allocating probability mass proportional to outcome rewards.
GFlowNets operate over a directed acyclic graph (DAG), where each path represents a sequence of actions from a root state to a terminal state. The flow through each path determines the probability of generating a particular outcome, enabling diverse and reward-aligned generation.
Formally, let $G=(\mathcal{S},\mathcal{A})$ denote the DAG, where nodes $s\in\mathcal{S}$ are states and directed edges $(s_{1}\to s_{2})\in\mathcal{A}$ are actions. If $(s_{1}\to s_{2})$ exists, then $s_{2}$ is a child of $s_{1}$, and $s_{1}$ is its parent. The graph has a unique initial state $s_{0}$ with no parents, and a set of terminal states $\mathcal{Y}$ with no children. A trajectory $\tau\in\mathcal{T}$ is a sequence of transitions $\tau=(s_{m}\to s_{m+1}\to\cdots\to s_{n})$.
A forward policy is a set of distributions $P_{F}(-|s)$ over the children of each nonterminal state $s\in\mathcal{S}$, inducing a trajectory distribution:
(2) $P_{F}(\tau=(s_{0}\to\cdots\to s_{n}))=\prod_{i=0}^{n-1}P_{F}(s_{i+1}\mid s_{i}).$
@@
2.4. Training Objective for GFlowNets
The sum over all $\tau\in\mathcal{T}$ in Eq. (3) is generally intractable, prompting the use of auxiliary variables to facilitate learning. We adopt the Subtrajectory Balance (SubTB) objective (Madan et al., 2023), which introduces the state flow $F(s)$ — a nonnegative scalar representing the total flow through state $s$:
(4) $F(s):=F(\{\tau\in\mathcal{T}:s\in\tau\})=\sum_{\tau\in\mathcal{T}:s\in\tau}F(\tau).$
Subtrajectory Balance. For any trajectory segment $\tau_{m:n}=(s_{m}\to\cdots\to s_{n})$, the following constraint holds:
@@ 2.4. Training Objective for GFlowNets
where $P_{F}(s_{i+1}|s_{i};\theta)$ is the forward policy, parameterized by $\theta$, representing the probability of taking action $(s_{i}\to s_{i+1})$ conditioned on the state $s_{i}$. Similarly, $P_{B}(s_{i}|s_{i+1};\theta)$ is the backward policy, modeling the reverse transitions. Here, $F(y)=R(y)$ if $y$ is terminal, and $F(s_{0})=Z$ if $s_{0}$ is the initial state.
@@ 3.1. Problem Formulation
Let $\mathcal{I}$ denote the set of valid items in the dataset, where each item $i\in\mathcal{I}$ is represented by its title—a token sequence $\mathbf{y}^{(i)}=[y_{1}^{(i)},y_{2}^{(i)},\cdots,y_{T}^{(i)}]$, with each token $y_{t}^{(i)}$ belonging to the LLM’s vocabulary. The titles collectively form a prefix tree, where shared prefixes define common paths. For instance, “Back to the Future II” and “Back to Life” share the prefix “Back to” before diverging. Note that in this example, words are used as tokens for illustration purposes. In practice, an item is automatically tokenized by an open-source LLM, such as LLaMA or Qwen.
This prefix tree is a special DAG: nodes represent tokens, and edges denote valid transitions to the next token. The root corresponds to the empty sequence, and each path to a terminal node forms a complete item title. We define:
  • State $s_{t}$ as a prefix sequence $s_{t}=y_{\leq t}=[y_{0},\cdots,y_{t}]$, representing a node’s path from the root.
  • Action $a_{t}$ as appending a valid token $y_{t+1}$, transitioning from the current state $s_{t}$ to a child state $s_{t+1}$.
Fig. 3(b) shows an example prefix tree for the 7 movie titles in Fig. 3(a). At the node “to”, the state is “Back to”, and valid actions lead to “Back to School”, “Back to Life”, or “Back to the”.
The traversal aligns with the GFlowNets framework, where the probability of generating an item is the product of transition probabilities along its path. For example, the probability of “Back to Life” is computed as the product over “Back” $\to$ “to” $\to$ “Life”.
Remark: This formulation casts generation as traversal through a prefix tree, ensuring only valid tokens are sampled at each step. It aligns model outputs with the dataset while leveraging GFlowNets to promote diversity and reward-proportional generation.